OpenStudy (anonymous):

Tanya has 4 red marbles, 6 blue marbles, and 8 green marbles. Which expression can be used to find the percent of Tanya's marbles that are blue? Answer Choice; A.(6)(100) ------- (4+8) B.(6)(100) ------- 8 C.(6)(100) ------- (4+6+8) D.(6)(100) -------- (4)(6)(8)

OpenStudy (anonymous):

% blue = number of blue * 100/ total Can you plug the numbers in yourself?

OpenStudy (anonymous):

@Chlorophyll , I need help :/

OpenStudy (apoorvk):

total no. of marbles = 4+6+8 = 18 no. of blue marbles = 6 \[percentage = \frac{\text{no. of of blue marbles}}{\text{total no. of marbles}} \text{x} 100\]

OpenStudy (anonymous):

6 blue marbles total = 4 red marbles, 6 blue marbles, and 8 green marbles

OpenStudy (anonymous):

@Chlorophyll , So the answer is C ?

OpenStudy (anonymous):

@Chlorophyll or D ?

OpenStudy (anonymous):

D has the denominator NOT TOTAL!

myininaya (myininaya):

total means you add

OpenStudy (anonymous):

6* 100/ total